A study was conducted to examine the relationship between variable X (independent variable) and Y (dependent variable). The Et Intercept X variable Coefficients 95.84 1.13 Standard Error 116.65 0.132 Stat 0.821 8.547 P-value 0.418 0.0000005
Answer #1

Estimated value = 95.84 + 1.13*191 = 311.67

Residual = actual - predicted = 382 - 311.67 = 70.33

Option B) 70.33 is the answer.